OpenStudy (anonymous):

If F(x) = 2x - 1, find the zero of F -1(x). -1 1/2 1

OpenStudy (anonymous):

replace \(x\) by 0 and see what you get

OpenStudy (anonymous):

let f(x)=y y=2x-1 this gives, x = (y+1)/2 If f(x)=y, then f^-1(y) = x So, f^-1(y) = (y+1)/2 => f^-1(x) = (x+1)/2

OpenStudy (anonymous):

I think he asked about the inverse function? @satellite73

OpenStudy (anonymous):

Its SHE @AbhimanyuPudi lol and is the answer gonna be -1/2x - 1

OpenStudy (anonymous):

Ohhhhh sorrrrrry...@minnie123

OpenStudy (anonymous):

Are you asking f(-1) or f^-1(x)

OpenStudy (anonymous):

find the zero of F ^-1(x).

OpenStudy (anonymous):

Okay f^-1(x) = (x+1)/2 So, 0 = (x+1)/2 0 = x+1 Then x = ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

No.... 0 = x+1 Subtract 1 on both sides 0-1 = x+1-1 -1 = x

OpenStudy (anonymous):

oh yeah don't know what i was thinking... :/ thanks

OpenStudy (anonymous):

welcome:)